OpenStudy (anonymous):

debbie anderson cuts a pice of wood meassuring 41/8 inches into two equal pieces. how long is each piece? the length of each piece is_in.

thomaster (thomaster):

\(\dfrac{41}{8}/2=\dfrac{41}{16}\)in you can see it this way: When you divide a fraction, you can multiply with the inverse of the fraction. So you flip it upside down. You can rewrite 2 as a fraction : \(\dfrac{2}{1}\) So \(\large\dfrac{\dfrac{41}{8}}{\dfrac{2}{1}}=\dfrac{41}{8}*\dfrac{1}{2}=\dfrac{41*1}{8*2}=\dfrac{41}{16}\)

OpenStudy (anonymous):

it 21/6

OpenStudy (anonymous):

it 2 1/6

thomaster (thomaster):

You can simplify it to \(2\dfrac{9}{16}\)